OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −1/2 ? A) y=x/2−2 B) y=−x/2−1 C) y=2x+1 D) y=−2x−1

OpenStudy (anonymous):

it is B, because if you put it in the equation y-y1=m(x-x1) y-2=-1/2(x+6) y-2=-1/2x-3 y=-1/2x-1